xx Timers

Timers

Erik Schnetter <schnetter@gmail.com>

August 07 2013

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11disablexcactusxtimerxset 10Scope: privatex01  BOOLEAN

11Description: Disable Cactus timer set

11                  10  x01  Default: no

11

11disablextimerxtrees 10Scope: privatex01  BOOLEAN

11Description: Disable timer trees

11                  10  x01  Default: no

11

11outputxprecision 10Scope: privatex01  INT

11Description: Number of decimal places to use in standard output for timer tree

11 Range 10 x01 Default: 1
111:*  10number of decimal places

11

11thresholdxpercentage10Scope: privatex01  REAL

11Description: The percentage of the root timer below which timers are omitted

11  Range  10  x01  Default: 1.0

110:*  10

11

11verbose  10Scope: privatex01  BOOLEAN

11Description: Output (debug) messages when a timer is started or stopped

11                  10  x01  Default: no

11

11xmlxclock  10Scope: privatex01  STRING

11Description: Which clock to use in the XML timer output file

11 Range 10 x01 Default: gettimeofday
11.*  10must be a legal clock name

11

11outxdir  10Scope: shared from IOx01STRING

11

7 Interfaces

General Implements: timers

Adds header:

CactusTimer.hh CactusTimerSet.hh Timer.hh TimerTree.hh

8 Schedule This section lists all the variables which are assigned storage by thorn Carpet/Timers. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

CCTKxSTARTUP x timerxstartup x prepare hierarchical timers x

11x11Before: 11driverxstartup
11 x 11Language:11c
11 x 11Type: 11function

11

CCTKxSHUTDOWN x timerxshutdown x prepare hierarchical timers x

11x11After: 11driverxshutdown
11 x 11Language:11c
11 x 11Type: 11function